# F-LMM: Grounding Frozen Large Multimodal Models

### v1

@@ 5 Conclusion
In this work, we studied the limitation of existing grounding LMMs, i.e., the loss of general world knowledge and instruction-following ability in the course of seeking state-of-the-art performance on grounding tasks. Regarding this issue, we make the first attempt to ground completely frozen LMMs that are well-trained for user-model conversation, based on the observation that geometric and spatial cues necessary for visual grounding already exist in LMMs’ self-attention mechanism. With the help of a CNN-based mask decoder and a SAM-based mask refiner, we achieve competitive visual grounding performance without losing any conversational ability of off-the-shelf LMMs. With both excellent conversational and visual grounding capabilities, the LMMs have the potential to perform complex visual perception and reasoning tasks like visual chain-of-thought reasoning.
6 Acknowledgements
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-PhD-2023-08-048T), the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Appendix A Appendix
In Sec A.1, we provide more detailed experimental results on both question-answering benchmarks and grounding benchmarks. Besides, we provide visualisation results in Sec A.3, including failure cases of existing grounding LMMs on general question-answering tasks, attention maps as well as segmentation masks, and examples of visual CoT and grounded conversations.
A.1 Benchmark results
Question-Answering Benchmarks. In addition to the four benchmarks reported in Table 1 of the main text, we also test the grounding LMMs on a wider range of question-answering benchmarks as shown in Table A1. Due to corrupted instruction-following abilities, existing grounding LMMs obtain zero or near-zero scores on these question-answering benchmarks.
@@ A.1 Benchmark results
Model MME MMB MMVet LLaVAW POPE GQA VQAv2 AI2D
@@
Referring Expression Segmentation. The results reported in Table 1 only include scores on the Val subsets of RefCOCO, RefCOCO+ and RefCOCOg. Here, we provide the grounding LMMs’ performances on all their subsets in Table A2. The metric used for Referring Expression Segmentation (RES) is cIoU.
@@
Panoptic Narrative Grounding. In Table 1 of the main text, we only report individual mask recalls on thing and stuff objects as well as the overall average recall. Here, we additionally report the mask recalls on singular and plural object nouns as shown in Table A3. As expected, segmenting plural nouns that refer to multiple object instances is more challenging for all the tested models.
@@ A.2 Mask Decoder
A.3 Visualisation
@@
General Multimodal Question-Answering. In Figure A2, we show some examples of grounding LMMs performing general question-answering tasks. When prompted to answer with single words (e.g., yes or no), existing grounding LMMs (GLaMM [57], LISA [28], and PixelLM [58]) usually fail to follow the user instructions. Besides, we also observe that the grounding LMMs tend to misunderstand the user’s questions as segmentation requests and reply mask tokens, e.g., ‘[SEG]’. Furthermore, these grounding LMMs fail to recognise the celebrity (Musk) and famous natural spot, exhibiting a worse grasp of world knowledge compared with a general-purpose LMM (e.g., LLaVA [39]). In contrast, F-LMM inherits the virtues of general-purpose LMMs (e.g., LLaVA-1.5 [37]) in instruction following and world knowledge comprehension, thanks to the ‘Frozen’ design philosophy.
Figure A3: Visualisations of K-Means results and segmentation masks. The attention weights (top-left) are mapped to 2D binary masks by the mask decoder (top-right) and then further optimised by the mask refiner (bottom-left). The samples are taken from the validation split of the PNG dataset [17]. The model used is DeepseekVL-1.3B [42].
From Attention Maps to Segmentation Masks. We visualise the geometric and spatial cues necessary for visual grounding in the word-image attention maps by applying KMeans clustering to the stacked attention maps that are collected from all transformer layers and attention heads. The attention maps of multiple-word objects are merged by element-wise average. As shown in Figure A3, we observe that the pixels of objects are roughly clustered together (top-left). With the CNN-Based mask decoder, the attention weights are mapped to 2D binary masks (top-right), which are then further optimised by the SAM-based mask refiner (bottom-left).
Figure A4: Visual Chain-of-Thought Reasoning. The model used is DeepseekVL-1.3B, and the samples are taken from the test set of VisCoT dataset [60]. The LMM is first prompted to think about the question-related object, which is then grounded by the mask head of F-LMM. The region of the question-related object is cropped and fed to the LMM to help answer the question.
@@
Figure A5: Visualisations of grounded human-AI conversations. The key phrases or words in the conversations can be precisely localised by the mask head of F-LMM. The LMM used is DeepseekVL-1.3B.
Visual Chain-of-Thought Reasoning. Figure A4 shows more examples of visual CoT by F-LMM. The model used in these examples is DeepseekVL-1.3B. When the LMM is prompted to answer ‘What object is the most relevant to the question?’, the mask head of F-LMM grounds the LMM’s answer about the relevant object by generating a segmentation mask, the bounding box of which is used to crop the object region from the original image. Then, the cropped image region is fed to the LMM to obtain the final answer. As shown in Figure A4, the Visual CoT empowered by the LMM’s grounding ability is helpful when the LMM needs to focus on question-related regions for visual perception and reasoning.
@@
Visualisation of Grounded Conversation. In Figure A5, we show some examples of grounding conversation by F-LMM. Our F-LMM maintains the LMMs’ original ability to follow the user’s instruction and understand unusual scenarios (e.g., the man ironing at the back of a taxi) while being able to localise the keywords and phrases during the conversations precisely. The model used in these examples is DeepseekVL-1.3B.
Appendix B Broader Impact
@@
Appendix C Limitations
While the proposed F-LMM approach demonstrates promising results in preserving conversational abilities while enhancing visual grounding, there are several key limitations that warrant consideration. (1) Inherited Biases and Limitations: As the F-LMM method is built upon frozen pre-trained LMMs, it inherits any biases or limitations present in the underlying models. These could include demographic biases, skewed knowledge representations, or other undesirable properties. (2) Limited Modality Scope: This work primarily focuses on vision-language multimodal interactions, without exploring other important modalities such as video, audio, and 3D point clouds. Expanding the scope to these additional modalities is a great direction to explore in the future. (3) Model Size Constraints: The experiments were restricted to LMMs up to 8 billion in parameter counts due to limited computing resources. Larger and more powerful models beyond this scale were not included. To address these limitations, future research could focus on mitigating biases, expanding the modality scope, and exploring larger-scale models.

### v2

@@ 5 Conclusion
In this work, we have studied the limitation of existing grounding LMMs, i.e., the loss of general world knowledge and instruction-following ability. To address this issue, we make the first attempt to ground fully frozen LMMs, which are already well-trained for user-model conversation, based on the insight that the geometric and spatial cues needed for visual grounding are inherently present within the self-attention mechanism of LMMs. By incorporating a CNN-based mask decoder and a SAM-based mask refiner, we achieve competitive visual grounding performance without sacrificing any conversational abilities of pre-trained LMMs. With this combination of strong conversational and visual grounding capabilities, these LMMs show promise for complex perception and reasoning tasks, such as segmentation with reasoning, grounded conversation generation, and visual chain-of-thought reasoning.
A1 Appendix
In Sec A1.1, we provide more detailed experimental results on both question-answering benchmarks and grounding benchmarks. Then we introduce the architecture of F-LMM’s mask decoder in Sec A1.2. We also summarise the datasets used by F-LMM and existing grounding LMMs in Sec A1.3. Last, we provide visualisation results in Sec A1.4, including failure cases of existing grounding LMMs on general question-answering tasks and examples of reasoning segmentation, grounded conversations and visual CoT. The broader impact and limitations of this work are elucidated in Sec A1.5 and Sec A1.6, respectively.
A1.1 Benchmark results
Question-Answering Benchmarks. In addition to the four benchmarks reported in the main text, we test the grounding LMMs on a wider range of question-answering benchmarks as shown in Table A1. Due to corrupted instruction-following abilities, existing grounding LMMs obtain zero or near-zero scores on these benchmarks.
@@
Model MME MMBench MMVet LLaVAW POPE GQA VQAv2 AI2D
@@
Referring Expression Segmentation. The results reported in the main text only include scores on the Val subsets of RefCOCO, RefCOCO+ and RefCOCOg. Here, we provide the grounding LMMs’ performances on all their subsets in Table A2. The metric used for Referring Expression Segmentation (RES) is cIoU.
@@
Panoptic Narrative Grounding. In the main text, we only report individual mask recalls on thing and stuff objects as well as the overall average recall. Here, we additionally report the mask recalls on singular and plural object nouns as shown in Table A3. As expected, segmenting plural nouns that refer to multiple object instances is more challenging for all the tested models.
@@
Table A4: Datasets used by F-LMM and existing grounding LMMs. COCOs stands for the COCO-Stuff [5] dataset and COCOp is for the COCO-Panoptic [25] dataset.
Datasets Language-Based Segmentation Standard Segmentation Grounded Conversation
RefCOCO RefCLE VG PNG Flickr COCOs ADE Mapillary VLPart COCOp Cityscapes OpenImage GranD MUSE GVC M3G2
F-LMM (Ours)
LISA [28]
Llava-G [80]
GLaMM [60]
GroundHog [84]
PixelLM [61]
LaSagnA [70]
A1.3 Dataset Usage
Training Data Comparison. In Table A4, we show the datasets used by existing grounding LMMs and our F-LMM. Existing methods conduct training on a wide range of standard segmentation datasets for excellent grounding ability and collect grounded conversation datasets to preserve chat ability. In contrast, F-LMM only need the RefCOCO and PNG datasets for segmentation capability, without needing additional grounded instruction-tuning datasets.
Training Data Format. For PNG dataset, an image narrative (e.g., ‘A hot air balloon is flying over the river’) is formatted as ‘User: Describe the image. Model: A hot air balloon is flying over the river.’ The coloured texts indicate the keywords for grounding, which are annotated in PNG’s training set. For RES dataset where each image is associated with multiple referring expressions (e.g. ‘The man in blue T-short’, ‘The girl who is smiling’). We convert the RES data to PNG format by concatenating the referring expressions into a single sentence: ‘User: Describe the image. Model: The man in blue T-short; The girl who is smiling.’
A1.4 Visualisation
@@
General Multimodal Question-Answering. In Figure A2, we show some examples of grounding LMMs performing general question-answering tasks. When prompted to answer with single words (e.g., yes or no), existing grounding LMMs (GLaMM [60], LISA [28], and PixelLM [61]) usually fail to follow the user instructions. Besides, we also observe that the grounding LMMs tend to misunderstand the user’s questions as segmentation requests and reply mask tokens, e.g., ‘[SEG]’. Furthermore, these grounding LMMs fail to recognise the celebrity (Musk) and famous natural spot, exhibiting a worse grasp of world knowledge compared with a general-purpose LMM (e.g., LLaVA [39]). In contrast, F-LMM inherits the virtues of general-purpose LMMs in instruction following and world knowledge comprehension, thanks to the ‘Frozen’ design philosophy.
Figure A3: Examples of Reasoning Segmentation. The red masks in the images are segmentation results. The model generating the answers is DeepSeekVL-7B. The sentences in the answers are grounded by the mask head of F-LMM.
@@
Figure A4: Visualisations of grounded human-AI conversations. The key phrases or words in the conversations can be precisely localised by the mask head of F-LMM. The LMM used is DeepSeekVL-7B.
Figure A5: Visual Chain-of-Thought Reasoning. The model used is DeepSeekVL-7B, and the samples are taken from the test set of VisCoT dataset [63]. The LMM is first prompted to think about the question-related object, which is then grounded by the mask head of F-LMM. The region of the question-related object is cropped and fed to the LMM to help answer the question.
Figure A6: Comparison between the attention maps of DeepSeekVL-7B(indicated as ‘D’) and LLaVA-1.5-7B(indicated as ‘L’). (m,n) means attention map at the n-th head of the m-th layer.
Reasoning Segmentation. We show examples of F-LMM performing reasoning segmentation in Figure A3. The LMM is first prompted to answer a question relevant to an object in the given image. The content of the answer is regarded as the grounding target. We observe that the LMM (DeepSeekVL-7B) is able to generate the correct answer of the queried object, which is then precisely localised by the mask head of F-LMM.
@@
Grounded Conversation. In Figure A4, we show some examples of grounding conversation by F-LMM. Our F-LMM maintains the LMMs’ original ability to follow the user’s instruction and understand unusual scenarios (e.g., the man ironing at the back of a taxi) while being able to localise the keywords and phrases during the conversations precisely. The model used in these examples is DeepSeekVL-7B.
Visual Chain-of-Thought Reasoning. Figure A5 shows examples of visual CoT by F-LMM. The model used in these examples is DeepSeekVL-7B. When the LMM is prompted to answer ‘which object is the most relevant to the question’, the mask head of F-LMM grounds the LMM’s answer about the relevant object by generating a segmentation mask, the bounding box of which is used to crop the object region from the original image. Then, the cropped image region is fed to the LMM to obtain the final answer. As shown in Figure A5, the Visual CoT empowered by the LMM’s grounding ability is helpful when the LMM needs to focus on question-related regions for visual perception and reasoning. We observe that the LMM can infer the location of text contents (e.g., fax number) in a document and the mask head can localise the relevant region, even though they are not trained on such data (i.e. VisCoT [63]).
Attention Maps of Different LMMs. In Figure A6, we compare the attention maps of DeepSeekVL-7B(indicated as ‘D’) and LLaVA-1.5-7B(indicated as ‘L’). (m,n) means attention map at the n-th head of the m-th layer. The two models are similar in showcasing shapes and locations of objects but differ in specific attention map patterns and textures. For example, LLaVA’s attention tends to be more concentrated with higher amplitudes.
A1.5 Broader Impact
@@
A1.6 Limitations
While the proposed F-LMM approach demonstrates promising results in preserving conversational abilities while enhancing visual grounding, there are several key limitations that warrant consideration.
  • Inherited Biases and Limitations: As the F-LMM method is built upon frozen pre-trained LMMs, it inherits any biases or limitations present in the underlying models. These could include demographic biases, skewed knowledge representations, or other undesirable properties.
  • Limited Modality Scope: This work primarily focuses on vision-language multimodal interactions, without exploring other important modalities such as video, audio, and 3D point clouds. Expanding the scope to these additional modalities is a great direction to explore in the future.
  • Model Size Constraints: The experiments were restricted to LMMs up to 8 billion in parameter counts due to limited computing resources. Larger and more powerful models beyond this scale were not included. To address these limitations, future research could focus on mitigating biases, expanding the modality scope, and exploring larger-scale models.
Acknowledgement. This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-PhD-2023-08-048T), the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).